OpenStudy (anonymous):

The graph of which quadratic equation has a vertex of (–2, 1)?

OpenStudy (luigi0210):

Do we need to make one up or..?

OpenStudy (anonymous):

information not enough

OpenStudy (anonymous):

that's why i'm confused i feel like the question doesn't have enough information here are the answers they give y = (x + 1)2 – 2 y = (x – 1)2 – 2 y = (x + 2)2 + 1 y = (x – 2)2 + 1 @Luigi0210 and @Jonask

OpenStudy (anonymous):

\[y=a(x-p)^2+q\]vertex (p,q) in normal cases you should have an extra point given however they aready showing u dat a=1

OpenStudy (anonymous):

so i just plug it in

OpenStudy (anonymous):

never mind i found out thank you

OpenStudy (anonymous):

yes the 3rd one is correct